# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v6

@@ 9.9 How to establish a benchmark?
As for the requirment of a wide range of scenes and tasks. Although iGibson (Li et al. (2021)) and BEHAVIOR-1K (Li et al. (2023a)) support simulating a variety of household tasks with high realism, they are still manually constructed. In Sec. 8, we discuss how foundation models can automate the generation of scenes and tasks. Using foundation models to create numerous scenes, combined with VLMs for accuracy checking and minimal human intervention, could be a valuable approach to explore.
@@ 10 Conclusion
Additionally, the proposed framework has certain limitations: (1) The framework is designed with a sequential structure, which contrasts with the parallel execution in human operation. (2) Both the proposed framework and the surveyed literature are based on learning-based approaches. While model-based methods may not generalize as well, they tend to significantly outperform learning-based methods in terms of success rates, precision and safety for specific tasks (Pang et al. (2023)). Therefore, investigating the integration of learning-based and model-based approaches remains an important research. (3) The framework proposed in this paper is based on the development of learning-based methods and the definition of general manipulation. The framework of brain-like cognitive research should also be explored.
@@
    Policy Current research on RFMs primarily involves fine-tuning VLMs. This approach deprives RFMs of the ability to self-explore (Wang et al. (2024b)). The extensive parameters of RFMs require significant computational resources for training and real-time reference, and model training also needs abundant data. Additionally, there is a lack of a unified benchmark for evaluating different RFMs.
@@
The authors thank the editors and anonymous reviewers for the constructive feedback and in-depth review of this work. This paper is dedicated to the researchers and engineers who aspire to advance the field of robotics — let’s change the world.
@@
As shown in Tab. 4 and Tab. 5. The types of datasets currently used can be categorized into internet image-language pairs data, human video data, and robot demonstration data in real-world and simulation environments. RT-2 (Brohan et al. (2023)) demonstrates that co-training with internet image-language pairs improves the model’s generalization. GR-1 (Wu et al. (2023a)), GR-2 (Cheang et al. (2024)), and LAPA (Ye et al. (2024)) show that training with human video data and robot demonstrations also enhances task generalization. GR-1 and GR-2 extract priors from human video data via image prediction, while LAPA uses latent-action for prior extraction. The advantages and disadvantages between these two methods are not clearly evident.
@@
The Input Modality shows that VLAKP primarily uses language and 3D representations as inputs, while VLADP relies on language, images, and proprioception. At the same time, adding a mask from the open-set visual module to the input improves the model’s generalization (Stone et al. (2023)). In terms of model architecture, VLAKP tends to use a 3D Feature Lift model combined with a Diffusion Model, while VLADP adopts a dual-system approach (slow and fast systems). This allows VLADP to effectively utilize the prior knowledge from VLM and enhance inference frequency.
@@
The current training objectives include next-token prediction, regression, diffusion, and TD learning in reinforcement learning. No research has conclusively shown which method is best. TD Learning in reinforcement learning can help robotic systems become more proficient than human teleoperators, exploiting the full potential of the hardware to perform tasks quickly, fluently, and reliably. It also enables robotic systems to improve autonomously through gathered experience, instead of relying solely on high-quality demonstrations. The Diffusion Objective can address action multi-modality issues and is widely applied in dexterous manipulation and high-dimensional action space tasks. Diffusion can be divided into Diffusion Policy-based and Flow Matching Policy-based approaches. Experiments in IMLE (Rana et al. (2025)) show that Diffusion Policy performs better than Flow Matching Policy. Fast (Pertsch et al. (2025)) introduces a compression-based tokenization scheme that matches the performance of diffusion VLAs. OpenVLA-OFT (Kim et al. (2025)) shows that the regression train-objective outperforms next-token prediction and diffusion. Therefore, different data, architectures, and tasks require different training objectives. The specific situations corresponding to each train objective should be deeply researched.
@@
From the Output Modality, it can be seen that outputting the hindsight goal image along with the action improves model stability (Bousmalis et al. (2023); Cheang et al. (2024)). The Benchmark shows that current benchmarks vary across methods. A unified benchmark standard can promote progress in the field. Regarding Success Rate, there are few experiments testing adaptation ability. Adaptation ability refers not only to quickly adapting to new tasks but also to quickly adapting to new embodiment configurations. Cross-embodiment learning can be divided into unified action space (Liu et al. (2024d); Black et al. ) and changing the action head (Team et al. (2024); Wang et al. (2024d)). Currently, the comparison between these two methods for cross-embodiment is not very clear. Regarding the manipulation task, it starts with single-arm pick-and-place tasks and evolves into dexterous manipulation using humanoid robot upper bodies. However, a considerable gap remains to achieve general dexterous manipulation. From the Failure Mode, it is observed that ”imprecise pose error” and ”wrong object error” occur most frequently.
@@
3D observation can be expressed in various forms. These include RGBD images, point clouds, voxels or multi-view images with camera extrinsic parameters (Ze et al. (2024b)). These 3D forms has not achieved large-scale adoption on the internet. As a result, the volume of 3D data remains significantly smaller than that of 2D images (Chen et al. (2024b)). Although SpatialVLM (Chen et al. (2024b)) uses off-the-shelf models to convert 2D images into 3D forms, the quality remains uncertain. Generating large-scale 3D observations in simulation environments might be an effective approach. However, there is still a sim-to-real gap.
@@
As for 2D and 3D representation learning, we have already introduced many pre-trained encoders for 2D representation in Sec. 4. For 3D representation encoders, the main options currently include PointNet++ (Qi et al. (2017)) and PointNext (Qian et al. (2022)). These encoders extract key features from point clouds. DP3 (Ze et al. (2024b)) introduces a holistic 1D embedding pooled from the 3D scene point cloud, which outperforms PointNet++ and PointNext. However, 3D Diffuser Actor (Ke et al. (2024)) shows that generating 3D representations by lifting features from perspective views to a 3D robot workspace, based on sensed depth and camera extrinsics, achieves even better results than DP3 (Ze et al. (2024b)). Additionally, converting 3D data into 2D allows the use of pre-trained 2D encoders that trained on large-scale datasets for feature extraction. Then, lifting 2D feature to 3D space. This method extracts texture features, spatial features and semantic features and become a notable trend in 3D representation.
@@
The current lift techniques can be categorized into three types: Direct Reconstruction, Feature Fusion, and Neural Field (Hong et al. (2023b)). Direct Reconstruction refers to features are mapped directly to the 3D space using camera extrinsics. However, this method is sensitive to noise in the camera pose. Feature Fusion combines 2D features into 3D maps using gradslam (Murthy Jatavallabhula et al. (2019)). This approach is more robust to camera pose noise. However, it requires depth map rendering from 3D data. Neural Field constructs 3D compact representation using a neural voxel field (Sun et al. (2022)). This method is more robust to noise in camera pose and does not require depth map renderings from 3D data.
@@
Current manipulation tasks are mainly divided into high-level and low-level. High-level tasks involve decision-making, such as the hierarchy of skills. Low-level tasks focus on execution, like policies. For high-level manipulation task, 3D observation has stronger spatial reasoning capabilities compared to 2D observation (Chen et al. (2024b)). It can recognize quantitative relationships of physical objects, such as distances or size differences. For low-level manipulation task, ChainedDiffuser (Xian et al. (2023)) demonstrates that 3D methods are more stable than 2D methods under varying camera viewpoints. DP3 (Ze et al. (2024b)) has shown that diffusion policies with 3D input achieve higher success rates compared to 2D image and the point cloud format performs best. However, the comparison is based on a relatively small dataset. Lin et al. (2024a) introduces a scaling law for object diversity and environment diversity. However, it is still unclear whether 2D or 3D observation is more suitable for the scaling law.
@@
Bicchi (2000) offers a thorough and widely accepted definition: dexterous manipulation is the capability of changing the position and orientation of the manipulated object from a given reference configuration to a different one, arbitrarily chosen within the hand workspace. Based on this definition, the dexterous manipulation can be described as: based on the designed end-effector, determining a sequence of contact points and the forces/torques to be exerted on the object, and control the whole-body to accomplish a specific task.
@@
As shown in Figure 9, we use a simple process to explain the generation of a sequence of contact points and forces/torques. Assuming the motion trajectory of the object is already obtained, the target wrench can be calculated based on the object’s mass and inertia. Contact points can be optimized based on object geometry, object material, and end-effector geometry etc, using appropriate metrics (Ferrari et al. (1992)). Alternatively, they can be generated using a knowledge-based approach (Stansfield (1991)). Subsequently, a coordinate system is established at the centroid of the object. Based on the location of the contact points, the target wrench is converted into fingertip force. Finally, the force of the fingertip is converted into the forces/torques required by the actuators through hand jacobian (Okamura et al. (2000)).
The pipeline contact points and forces/torques mentioned above are derived sequentially. In practical applications, both can be optimized simultaneously (Xu et al. (2024)). However, due to the discontinuities and mode changes inherent in contact dynamics, it creates challenges for planning in contact-rich manipulation (Pang (2023)).
As shown in Fig. 10, $q^{a}$ and $\mu$ represent the actual and commanded positions of the actuated ball, while $q^{u}$ represents an unactuated box that slides along a rail with sufficient damping. Our goal is to push the box by the ball to transition from the current configuration $q^{u}$ to the goal configuration $q^{u}_{goal}$. Suppose $q^{u}_{+}$ is the steady-state position of the box after the ball is commanded to position $\mu$. We expect that $q^{u}_{+}$ and $q^{u}_{goal}$ are sufficiently close. Typically, $q^{u}_{+}=f(q^{u},q^{a},\mu)$ represents the system’s dynamic, which can be obtained through system identification (Suh et al. (2023)).
@@
As shown in Fig. 10, the gradient is non-smooth when the ball is either in contact with or separated from the box. This presents challenges for optimization-based problem-solving methods. Current approaches for smoothing contact dynamics are mainly divided into analytic smoothing and randomized smoothing (Pang et al. (2023)). However, it is often necessary to consider frictional contact constraints, which can make the optimization problem non-convex (Tedrake (2023); Le Lidec et al. (2024)). To address this issue, current methods primarily rely on Anitescu’s convex relaxation(Anitescu (2006); Pang et al. (2023); Jin (2024)) and Todorov’s model (Todorov et al. (2012); Todorov (2014)). At the same time, the contact dynamics also need to consider the robot’s joint limits constraint, the penetration constraint between the hand and the object, and the self-penetration constraint (Pang and Tedrake (2021); Xu et al. (2024)). Although smoothed dynamics on trajectory optimization, such as iMPC (Suh et al. (2022) and LQR (Shirai et al. (2024)) are considered effective, these methods are only suitable when the goal is close to the initial configuration. When the goal is farther from the initial configuration, a more global strategy is required.
@@
As shown in Fig. 10, there are two initial guesses: 1) the ball is on the left of the box, and 2) the ball is on the right of the box. Different initial guesses lead to different dynamic model. Choosing the wrong initial guess can prevent the optimization problem finding the optimal solution. To address this issue, the system dynamics need to be divided into contact modes, and a global strategy is required (Pang et al. (2023)). In Fig. 10, the ball and the box experience both contact and no contact. However, for a system with $n$ rigid bodies, there can be up to $C^{2}_{n}$ contact pairs. Each contact pair may have multiple contact modes, such as sticking/rolling, separation, and sliding, which can lead to an explosion of contact modes. Although Huang et al. (2021) considers kinematically feasible contact modes and reduces the number to $O(n^{d^{2}/2+2.5d})$(where d is the number of degrees of freedom and n is the number of contact points), the reduced number of contact modes is still too large to handle.
@@
To address the above issues, current model-based techniques are mainly divided into contact-explicit and contact-implicit (Jiang et al. (2024b)). The contact-explicit approach in dexterous in-hand manipulation involves explicitly representing contacts, such as their locations, modes, and reaction forces (Cheng et al. (2023)). These representations form contact sequences, which are obtained through methods like enumerating (Aceituno-Cabezas and Rodriguez (2020); Hogan and Rodriguez (2020)), searching (Chen et al. (2021a); Zhu et al. (2023a); Cruciani et al. (2018)), sampling (Cheng et al. (2023)), or demonstration (Khadivar and Billard (2023)). However, these methods suffer from poor scalability (Pang et al. (2023)) and contacts easily fall into the local optimum (Jiang et al. (2024b)). The contact-implicit approach avoids the complexity of explicit contact representation. It uses methods like relaxed complementary constraints (Kim et al. (2023); Le Cleac’h et al. (2024)), smooth surrogate models (Pang et al. (2023); Önol et al. (2019)), or direct control sampling (Howell et al. (2022)). However, the smoothing process introduces force-at-a-distance effects and sacrifices physical fidelity (Pang et al. (2023)). Although Jiang et al. (2024b) compensates for discrepancies between actual and planned contact modes using tactile feedback, this method assumes quasi-static contact model and struggles to handle highly dynamic actions, such as spinning a pen between fingers.
@@
Table 4: Representative Policy Works in Sec. 7. The policy type is consistent with VLAC, VLAKP, VLADP in Sec. 7. Current robot foundation model training recipes include From-Scratch and Pre-Train + Fine-Tune (Post-Train). Some methods use vision or language foundation models during training. However, these foundation models are not pre-trained on large manipulation datasets. So these methods are still considered From-Scratch( Shridhar et al. (2021)). Pre-Train refers to training on extensive manipulation datasets. Fine-Tune involves post-training a pre-trained model on a small task-specific manipulation dataset for a particular task. The datasets used in various stages are detailed in the ‘Dataset’. ‘Benchmark’ refers to the benchmark used in the experiment. The ‘Success Rate’ is categorized into Seen and Unseen. ‘Seen’ refers to test cases that have appeared in the training data, while ‘Unseen’ refers to test cases that have not been present in the training data. However, some studies do not clearly differentiate between Seen and Unseen cases, and in our analysis, we also do not make this distinction. Meanwhile, for the pre-train + fine-tune training recipe, the success rate is further divided into the out-of-box generalization ability of the pre-trained model and the adaptation ability of the fine-tuned model. We have also provided annotations on the success rates of both the out-of-box and fine-tuned models. Due to differences in benchmarks, solely estimating the success rate is meaningless. Therefore, we also include the success rates of the baseline for comparison. ‘Embodiment Config’ refers to the type of robot selected in the experiment. ‘Manipulation Task’ refers to the manipulation task designed for the experiment. ‘Failure Mode’ refers to the common failure cases of the method, as indicated in the paper.
@@
Code as Policy (Liang et al. (2023))
@@
VoxPoser (Huang et al. (2023c)) VLAKP NA NA NA Self-Creation
@@
ReKep (Huang et al. (2024a)) VLAKP NA NA NA Self-Creation
@@
CLIPort (Shridhar et al. (2021)) VLAKP
@@
PerAct (Shridhar et al. (2023)) VLAKP
@@
ChainedDiffuser (Xian et al. (2023)) VLAKP
@@
3D Diffuser Actor (Ke et al. (2024)) VLAKP
@@
ManiFoundation ((Huang et al. (2023a))) VLAKP NA
@@
Gato (Reed et al. (2022)) VLADP NA
@@
RoboCat (Bousmalis et al. (2023)) VLADP NA
@@
RoboAgent (Bharadhwaj et al. (2023)) VLADP NA

### v7

@@ 9.9 How to establish a benchmark?
As for the requirment of a wide range of scenes and tasks. Although iGibson (li2021igibson) and BEHAVIOR-1K (li2023behavior) support simulating a variety of household tasks with high realism, they are still manually constructed. In Sec. 8, we discuss how foundation models can automate the generation of scenes and tasks. Using foundation models to create numerous scenes, combined with VLMs for accuracy checking and minimal human intervention, could be a valuable approach to explore.
@@ 10 Conclusion
Additionally, the proposed framework has certain limitations: (1) The framework is designed with a sequential structure, which contrasts with the parallel execution in human operation. (2) Both the proposed framework and the surveyed literature are based on learning-based approaches. While model-based methods may not generalize as well, they tend to significantly outperform learning-based methods in terms of success rates, precision and safety for specific tasks (pang2023global). Therefore, investigating the integration of learning-based and model-based approaches remains an important research. (3) The framework proposed in this paper is based on the development of learning-based methods and the definition of general manipulation. The framework of brain-like cognitive research should also be explored.
@@
    Policy Current research on RFMs primarily involves fine-tuning VLMs. This approach deprives RFMs of the ability to self-explore. The extensive parameters of RFMs require significant computational resources for training and real-time reference, and model training also needs abundant data. Additionally, there is a lack of a unified benchmark for evaluating different RFMs.
@@
The authors thank the editors and anonymous reviewers for the constructive feedback and in-depth review of this work. This paper pays tribute to the researchers and engineers who tirelessly advance the field of robotics — let’s change the world.
@@
As shown in Tab. 4 and Tab. 5. The types of datasets currently used can be categorized into internet image-language pairs data, human video data, and robot demonstration data in real-world and simulation environments. RT-2 (rt2) demonstrates that co-training with internet image-language pairs improves the model’s generalization. GR-1 (wu2023unleashing), GR-2 (cheang2024gr), and LAPA (ye2024latent) show that training with human video data and robot demonstrations also enhances task generalization. GR-1 and GR-2 extract priors from human video data via image prediction, while LAPA uses latent-action for prior extraction. The advantages and disadvantages between these two methods are not clearly evident.
@@
The Input Modality shows that VLAKP primarily uses language and 3D representations as inputs, while VLADP relies on language, images, and proprioception. At the same time, adding a mask from the open-set visual module to the input improves the model’s generalization (stone2023open). In terms of model architecture, VLAKP tends to use a 3D Feature Lift model combined with a Diffusion Model, while VLADP adopts a dual-system approach (slow and fast systems). This allows VLADP to effectively utilize the prior knowledge from VLM and enhance inference frequency.
@@
The current training objectives include next-token prediction, regression, diffusion, and TD learning in reinforcement learning. No research has conclusively shown which method is best. TD Learning in reinforcement learning can help robotic systems become more proficient than human teleoperators, exploiting the full potential of the hardware to perform tasks quickly, fluently, and reliably. It also enables robotic systems to improve autonomously through gathered experience, instead of relying solely on high-quality demonstrations. The Diffusion Objective can address action multi-modality issues and is widely applied in dexterous manipulation and high-dimensional action space tasks. Diffusion can be divided into Diffusion Policy-based and Flow Matching Policy-based approaches. Experiments in IMLE (rana2025imle) show that Diffusion Policy performs better than Flow Matching Policy. Fast (pertsch2025fast) introduces a compression-based tokenization scheme that matches the performance of diffusion VLAs. OpenVLA-OFT (kim2025fine) shows that the regression train-objective outperforms next-token prediction and diffusion. Therefore, different data, architectures, and tasks require different training objectives. The specific situations corresponding to each train objective should be deeply researched.
@@
From the Output Modality, it can be seen that outputting the hindsight goal image along with the action improves model stability (bousmalis2023robocat; cheang2024gr). The Benchmark shows that current benchmarks vary across methods. A unified benchmark standard can promote progress in the field. Regarding Success Rate, there are few experiments testing adaptation ability. Adaptation ability refers not only to quickly adapting to new tasks but also to quickly adapting to new embodiment configurations. Cross-embodiment learning can be divided into unified action space (liu2024rdt; black2410pi0) and changing the action head (team2024octo; wang2024scaling). Currently, the comparison between these two methods for cross-embodiment is not very clear. Regarding the manipulation task, it starts with single-arm pick-and-place tasks and evolves into dexterous manipulation using humanoid robot upper bodies. However, a considerable gap remains to achieve general dexterous manipulation. From the Failure Mode, it is observed that ”imprecise pose error” and ”wrong object error” occur most frequently.
@@
3D observation can be expressed in various forms. These include RGBD images, point clouds, voxels or multi-view images with camera extrinsic parameters (ze20243d). These 3D forms has not achieved large-scale adoption on the internet. As a result, the volume of 3D data remains significantly smaller than that of 2D images (chen2024spatialvlm). Although SpatialVLM (chen2024spatialvlm) uses off-the-shelf models to convert 2D images into 3D forms, the quality remains uncertain. Generating large-scale 3D observations in simulation environments might be an effective approach. However, there is still a sim-to-real gap.
@@
As for 2D and 3D representation learning, we have already introduced many pre-trained encoders for 2D representation in Sec. 4. For 3D representation encoders, the main options currently include PointNet++ (qi2017pointnet++) and PointNext (qian2022pointnext). These encoders extract key features from point clouds. DP3 (ze20243d) introduces a holistic 1D embedding pooled from the 3D scene point cloud, which outperforms PointNet++ and PointNext. However, 3D Diffuser Actor (ke20243d) shows that generating 3D representations by lifting features from perspective views to a 3D robot workspace, based on sensed depth and camera extrinsics, achieves even better results than DP3 (ze20243d). Additionally, converting 3D data into 2D allows the use of pre-trained 2D encoders that trained on large-scale datasets for feature extraction. Then, lifting 2D feature to 3D space. This method extracts texture features, spatial features and semantic features and become a notable trend in 3D representation.
@@
The current lift techniques can be categorized into three types: Direct Reconstruction, Feature Fusion, and Neural Field (hong20233d). Direct Reconstruction refers to features are mapped directly to the 3D space using camera extrinsics. However, this method is sensitive to noise in the camera pose. Feature Fusion combines 2D features into 3D maps using gradslam (murthy2019gradslam). This approach is more robust to camera pose noise. However, it requires depth map rendering from 3D data. Neural Field constructs 3D compact representation using a neural voxel field (sun2022direct). This method is more robust to noise in camera pose and does not require depth map renderings from 3D data.
@@
Current manipulation tasks are mainly divided into high-level and low-level. High-level tasks involve decision-making, such as the hierarchy of skills. Low-level tasks focus on execution, like policies. For high-level manipulation task, 3D observation has stronger spatial reasoning capabilities compared to 2D observation (chen2024spatialvlm). It can recognize quantitative relationships of physical objects, such as distances or size differences. For low-level manipulation task, ChainedDiffuser (xian2023chaineddiffuser) demonstrates that 3D methods are more stable than 2D methods under varying camera viewpoints. DP3 (ze20243d) has shown that diffusion policies with 3D input achieve higher success rates compared to 2D image and the point cloud format performs best. However, the comparison is based on a relatively small dataset. lin2024data introduces a scaling law for object diversity and environment diversity. However, it is still unclear whether 2D or 3D observation is more suitable for the scaling law.
@@
bicchi2000hands offers a thorough and widely accepted definition: dexterous manipulation is the capability of changing the position and orientation of the manipulated object from a given reference configuration to a different one, arbitrarily chosen within the hand workspace. Based on this definition, the dexterous manipulation can be described as: based on the designed end-effector, determining a sequence of contact points and the forces/torques to be exerted on the object, and control the whole-body to accomplish a specific task.
@@
As shown in Figure 9, we use a simple process to explain the generation of a sequence of contact points and forces/torques. Assuming the motion trajectory of the object is already obtained, the target wrench can be calculated based on the object’s mass and inertia. Contact points can be optimized based on object geometry, object material, and end-effector geometry etc, using appropriate metrics (ferrari1992planning). Alternatively, they can be generated using a knowledge-based approach (stansfield1991robotic). Subsequently, a coordinate system is established at the centroid of the object. Based on the location of the contact points, the target wrench is converted into fingertip force. Finally, the force of the fingertip is converted into the forces/torques required by the actuators through hand jacobian (okamura2000overview).
The pipeline contact points and forces/torques mentioned above are derived sequentially. In practical applications, both can be optimized simultaneously (xu2024manifoundation). However, due to the discontinuities and mode changes inherent in contact dynamics, it creates challenges for planning in contact-rich manipulation (pang2023planning).
As shown in Fig. 10, $q^{a}$ and $\mu$ represent the actual and commanded positions of the actuated ball, while $q^{u}$ represents an unactuated box that slides along a rail with sufficient damping. Our goal is to push the box by the ball to transition from the current configuration $q^{u}$ to the goal configuration $q^{u}_{goal}$. Suppose $q^{u}_{+}$ is the steady-state position of the box after the ball is commanded to position $\mu$. We expect that $q^{u}_{+}$ and $q^{u}_{goal}$ are sufficiently close. Typically, $q^{u}_{+}=f(q^{u},q^{a},\mu)$ represents the system’s dynamic, which can be obtained through system identification (suhdoes).
@@
As shown in Fig. 10, the gradient is non-smooth when the ball is either in contact with or separated from the box. This presents challenges for optimization-based problem-solving methods. Current approaches for smoothing contact dynamics are mainly divided into analytic smoothing and randomized smoothing (pang2023global). However, it is often necessary to consider frictional contact constraints, which can make the optimization problem non-convex (tedrake5underactuated; le2024contact). To address this issue, current methods primarily rely on Anitescu’s convex relaxation(anitescu2006optimization; pang2023global; jin2024complementarity) and Todorov’s model (todorov2012mujoco; todorov2014convex). At the same time, the contact dynamics also need to consider the robot’s joint limits constraint, the penetration constraint between the hand and the object, and the self-penetration constraint (pang2021convex; xu2024manifoundation). Although smoothed dynamics on trajectory optimization, such as iMPC (suh2022bundled and LQR (shirai2024linear) are considered effective, these methods are only suitable when the goal is close to the initial configuration. When the goal is farther from the initial configuration, a more global strategy is required.
@@
As shown in Fig. 10, there are two initial guesses: 1) the ball is on the left of the box, and 2) the ball is on the right of the box. Different initial guesses lead to different dynamic model. Choosing the wrong initial guess can prevent the optimization problem finding the optimal solution. To address this issue, the system dynamics need to be divided into contact modes, and a global strategy is required (pang2023global). In Fig. 10, the ball and the box experience both contact and no contact. However, for a system with $n$ rigid bodies, there can be up to $C^{2}_{n}$ contact pairs. Each contact pair may have multiple contact modes, such as sticking/rolling, separation, and sliding, which can lead to an explosion of contact modes. Although huang2021efficient considers kinematically feasible contact modes and reduces the number to $O(n^{d^{2}/2+2.5d})$(where d is the number of degrees of freedom and n is the number of contact points), the reduced number of contact modes is still too large to handle.
@@
To address the above issues, current model-based techniques are mainly divided into contact-explicit and contact-implicit (jiang2024contact). The contact-explicit approach in dexterous in-hand manipulation involves explicitly representing contacts, such as their locations, modes, and reaction forces (cheng2023enhancing). These representations form contact sequences, which are obtained through methods like enumerating (aceituno2020global; hogan2020feedback), searching (chen2021trajectotree; zhu2023efficient; cruciani2018dexterous), sampling (cheng2023enhancing), or demonstration (khadivar2023adaptive). However, these methods suffer from poor scalability (pang2023global) and contacts easily fall into the local optimum (jiang2024contact). The contact-implicit approach avoids the complexity of explicit contact representation. It uses methods like relaxed complementary constraints (kim2023contact; le2024fast), smooth surrogate models (pang2023global; onol2019contact), or direct control sampling (howell2022predictive). However, the smoothing process introduces force-at-a-distance effects and sacrifices physical fidelity (pang2023global). Although jiang2024contact compensates for discrepancies between actual and planned contact modes using tactile feedback, this method assumes quasi-static contact model and struggles to handle highly dynamic actions, such as spinning a pen between fingers.
@@
Table 4: Representative Policy Works in Sec. 7. The policy type is consistent with VLAC, VLAKP, VLADP in Sec. 7. Current robot foundation model training recipes include From-Scratch and Pre-Train + Fine-Tune (Post-Train). Some methods use vision or language foundation models during training. However, these foundation models are not pre-trained on large manipulation datasets. So these methods are still considered From-Scratch( shridhar2021cliport). Pre-Train refers to training on extensive manipulation datasets. Fine-Tune involves post-training a pre-trained model on a small task-specific manipulation dataset for a particular task. The datasets used in various stages are detailed in the ‘Dataset’. ‘Benchmark’ refers to the benchmark used in the experiment. The ‘Success Rate’ is categorized into Seen and Unseen. ‘Seen’ refers to test cases that have appeared in the training data, while ‘Unseen’ refers to test cases that have not been present in the training data. However, some studies do not clearly differentiate between Seen and Unseen cases, and in our analysis, we also do not make this distinction. Meanwhile, for the pre-train + fine-tune training recipe, the success rate is further divided into the out-of-box generalization ability of the pre-trained model and the adaptation ability of the fine-tuned model. We have also provided annotations on the success rates of both the out-of-box and fine-tuned models. Due to differences in benchmarks, solely estimating the success rate is meaningless. Therefore, we also include the success rates of the baseline for comparison. ‘Embodiment Config’ refers to the type of robot selected in the experiment. ‘Manipulation Task’ refers to the manipulation task designed for the experiment. ‘Failure Mode’ refers to the common failure cases of the method, as indicated in the paper.
@@
Code as Policy (liang2023code)
@@
VoxPoser (huang2023voxposer) VLAKP NA NA NA Self-Creation
@@
ReKep (huang2024rekep) VLAKP NA NA NA Self-Creation
@@
CLIPort (shridhar2021cliport) VLAKP
@@
PerAct (shridhar2023perceiver) VLAKP
@@
ChainedDiffuser (xian2023chaineddiffuser) VLAKP
@@
3D Diffuser Actor (ke20243d) VLAKP
@@
ManiFoundation ((huang2023embodied)) VLAKP NA
@@
Gato (reed2022generalist) VLADP NA
@@
RoboCat (bousmalis2023robocat) VLADP NA
@@
RoboAgent (bharadhwaj2023roboagent) VLADP NA